OpenStudy (anonymous):

Let f : R → Z, where f(x) = 2x − 1. (a) Find f(A), where A = {x ∈ R | 1 ≤ x ≤ 4}. (b) Find f ^-1 where B = {−9, −8}.

OpenStudy (amistre64):

so you have a line, and you want to know what values are for all xs from 1 to 4,

OpenStudy (amistre64):

well all integer values of f(A) that is

OpenStudy (amistre64):

one approach is to simply start converting x into f(x) 1 < x < 4 1(2) < 2x < 4(2) 1(2) -1 < 2x-1 < 4(2) -1

OpenStudy (anonymous):

do i really just need to plug it in? This is for discrete math. seems way too simple O.o

OpenStudy (anonymous):

Its for our part on Set theory, it really didn't occur to me just to plug it in.

OpenStudy (amistre64):

the problem is simple just so that you would be able to determine if your solution is correct i spose.

OpenStudy (anonymous):

Why explain that Real numbers map onto integers? any reason?

OpenStudy (amistre64):

domain (x inputs) are real values the range f(x) is composed of integers, so its prolly a stepping function

OpenStudy (amistre64):

other than that, its just information that you use to determine the required outcome

OpenStudy (anonymous):

However, when f is inverted, wouldn't it create some constraint as it makes a fraction for the set of integers? Or I'm just reading to far into it?

OpenStudy (amistre64):

\[\{f(A) \in Z~|~1\le f(A)\le7\} \] or some such stuff

OpenStudy (amistre64):

y = 2x-1 (y+1)/2 when y=1,2,3,4,5,6,7 (2,3,4,5,6,7,8)/2 are real values

OpenStudy (perl):

Assuming Z stands for set of integers, it looks like non integer outputs are undefined. f(2.3) is therefore undefined, since it is equal to 3.6 which is not an integer

OpenStudy (amistre64):

then if memory serves. f does not map real to integer if that is the case

OpenStudy (anonymous):

exactly so if I plug in -9 I would get an output of -4, which would be valid but the opposite would occur for -8. And yes, it does stand for set of integers

OpenStudy (perl):

but f(2.5) is defined

OpenStudy (amistre64):

which is why i was thinking of a step function

OpenStudy (amistre64):

the domain of the inverse is Z,

OpenStudy (anonymous):

Well, by definition ( what I've been taught), A ---> B would mean A means that B is mapped onto A* Im sorry

OpenStudy (anonymous):

means that B is mapped onto A *

OpenStudy (amistre64):

R to Z is function notation for inputs of R and outputs of Z, not sure why your definition would have it any other way

OpenStudy (anonymous):

Well, the books and teacher explanations are different ( maybe?), I mean I don't want to argue about it, that was in the book. By the teacher it said that Z would be the codomain and R would be the domain

OpenStudy (amistre64):

 granted that the solution set is R,Z is a set if discrete points.